Please help.
Subtract and write answer in simplest form.
y-2/y+3 - y/2y+6
I am getting y-4/2(y+3)
Is this right?

Answer by fractalier(6550)

From
y-2/y+3 - y/2y+6, we can multiply the first fractaion by 2/2 in order to combine it with the second fraction...so we get
y-2/y+3 - y/2y+6 =
(2y - 4) / (2y + 6) - y / (2y + 6) =
(y - 4) / (2y + 6) or
(y - 4) / 2(y + 3)
Looks like what you got! Way to go!